A table is three times as long as it is wide. If it were 3 ft shorter and 3 ft wider, it would be square (with all sides equal). How long and how wide is the table?
Answer by ankor@dixie-net.com(22740)

A table is three times as long as it is wide. If it were 3 ft shorter and 3 ft wider, it would be square (with all sides equal). How long and how wide is the table?
:
Let x = original width
then
3x = original length
:
If length is increased by 3 and the width is decreased by 3,it is square, therefore:
(3x - 3) = (x + 3)
3x -x = 3 + 3
2x = 6
x = 3 is the width, therefore length: 3*3 = 9